OpenStudy (anonymous):

Would someone help please? Thank you, Ill give a medal. Nancy left a bin outside in her garden to collect rainwater. She notices that 1 over 2 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words. I honestly just dont understand this question,

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

means that 1/2 {half} gallon is enough to fill 2/3{two third} of the bin so let x be the quantity of water in bin lets say x so 1-2/3=1/3 1/2g=2/3b 1/2g + xg =2/3b + 1/3b 1/2g+ xg= 1 further more its not coming yo me